OpenStudy (anonymous):

the sum of five times a number and three is equal to the difference of the number and thirteen. find the number.

OpenStudy (anonymous):

@boyb39 @kc_kennylau

OpenStudy (kc_kennylau):

5n+3=n-13

OpenStudy (anonymous):

what do I do now?

OpenStudy (kc_kennylau):

Solve that equation :)

OpenStudy (anonymous):

how...

OpenStudy (kc_kennylau):

Subtract n from both sides, subtract 3 from both sides

OpenStudy (anonymous):

and how would that look -5n=10?

OpenStudy (kc_kennylau):

how did you get -5n

OpenStudy (kc_kennylau):

5n+3=n-13 5n+3-n=n-13-n (subtract n from both sides) 4n+3=-13 4n+3-3=-13-3 (subtract 3 from both sides) 4n=-16

OpenStudy (anonymous):

n=-4?

OpenStudy (kc_kennylau):

yep